OpenStudy (peterg2001):

help will medal Add. Simplify the answer to lowest terms before entering the numerator and denominator in their boxes. 47/148 + 29/148

OpenStudy (peterg2001):

@dtuniyants

OpenStudy (chosenmatt):

add all the top numbers and then reduce it the denominator stays the same:)

OpenStudy (anonymous):

you add them together so 47 + 29 = 76 / 148 and simplify

OpenStudy (anonymous):

the same thing, just add the numenators and symplify: 47+29=76, so 76/148. divide 76 and 148 by 2: 38/74, divide again:19/37

OpenStudy (yanasidlinskiy):

Since denominators are common add and simplify to the simplest form if needed:)

OpenStudy (chosenmatt):

yup...

OpenStudy (anonymous):

then 76/148 you devide by 4 and get 19 / 37 l

OpenStudy (anonymous):

@YanaSidlinskiy